Domenica buys a 0.056 kilogram segment of her favourite blue vien cheese at a cost o $25.42 per kilogram. How much does this piece of cheese cost? Round off to the nearest 5 cents
Answer by ptaylor(2198)

Let x=cost of the segment of cheese
We can now set up a simple ratio:
1kg is to $25.42 as 0.056kg is to x and this can be written as:
1/(25.42)=0.056/x multiply both sides by (25.42x) to get rid of fractions
1(25.42x)/(25.42)=0.056(25.42x)/x simplify
x=0.056(25.42)=$1.4235~~~$1.40----------------for the segment of cheese
Check
How many (56/1000th) are there in (1000/1000th or 1)? Let k=that number
(56/1000)k=(1000/1000) multiply both sides by 1000
k=17.8571
(17.8571)($1.4235)=$25.42
~$25.42=$25.42

Hope this helps---ptaylor